# Rigorous Test for Quantum Integrability and Nonintegrability

### v2

@@ III.1 String Diagram
Before proving Thm. 2, we introduce a “basis-independent” approach for a proof of nonintegrability. We adopt a string diagram as a useful graphical language, which is widely used in the various fields of physics and mathematics [55, 56, 57]. The basic idea is to represent a linear map by a box and wires. For example, $\hat{X}\in\mathcal{B}\otimes\mathcal{B}_{0}$, $f:\mathcal{B}_{0}^{\otimes 3}\to\mathcal{B}_{0}^{\otimes 2}$, and $(\mathrm{id}_{\mathcal{B}}\otimes f)(\hat{X}\otimes\bullet):\mathcal{B}_{0}^{\otimes 2}\to\mathcal{B}\otimes\mathcal{B}_{0}^{\otimes 2}$ are denoted as
@@
respectively, where $\mathrm{id}_{\mathcal{B}}$ is the identity map on $\mathcal{B}$. Here, each box represents a linear map and the top line corresponds to an input and the bottom line corresponds to an output. In other words, time flows from top to bottom. Note that an element of a linear space is identified with a linear map from $\mathbb{C}$, which is denoted as the “empty” line. The horizontal direction corresponds to the spatial direction, and placing the boxes horizontally corresponds to their tensor product. In the following, solid and dashed lines represent $\mathcal{B}_{0}$ and $\mathcal{B}$, respectively. Using this expression, the injectivity of $[\hat{H}^{(2)}_{i},\bullet\otimes I]$ and $[\hat{H}^{(2)}_{i},I\otimes\bullet]$ (see Eq. (4)) is equivalent to the existence of linear maps $\mu_{i},\lambda_{i}:\mathcal{B}_{0}^{\otimes 2}\to\mathcal{B}_{0}$ satisfying
@@
III.2 Isomorphisms between $\mathcal{B}^{(k)}_{\leq}$’s
Now we move on to the proof of Eq. (7). It is sufficient to show the following proposition.
@@ III.2 Isomorphisms between $\mathcal{B}^{(k)}_{\leq}$’s
for all $i\in\Lambda_{N}$, which is equivalent to the condition $\hat{Q}_{[k-1]}^{(k-1)}\in\mathcal{B}_{\leq}^{(k-1)}$. The uniqueness of $\hat{Q}_{[k-1]}^{(k-1)}$ is also shown by the injectivity (23). ∎
If we define $\hat{Q}_{[k]}^{(k)}\in\mathcal{B}^{(k)}$ as in Eq. (24) for some $\hat{Q}_{[k-1]}^{(k-1)}\in\mathcal{B}_{\leq}^{(k-1)}$, then it is easy to verify $\hat{Q}_{[k]}^{(k)}\in\mathcal{B}_{\leq}^{(k)}$. Therefore, we have shown $\mathcal{B}_{\leq}^{(k)}\cong\mathcal{B}_{\leq}^{(k-1)}$ and Eq. (7). By applying Eq. (24) iteratively, $\hat{Q}^{(k)}_{[k],i}$ is represented as
@@
where $\hat{Q}_{[2]}^{(2)}=\sum_{i=1}^{N}\hat{H}^{(2)}_{[2],i}$ is an element of $\mathcal{B}^{(2)}_{\leq}$. We note that, if $\hat{Q}_{[2]}^{(2)}=\hat{H}^{(2)}$, this operator has the same form as the $3$-local conserved quantities of Bethe solvable fundamental systems [34] (see Eq. (10)).
III.3 Isomorphisms between <$\mathcal{B}^{(k)}_{<}$’s
We conclude this section by proving Eq. (8). Here, we shall prove only the case of $\mathcal{B}_{<}^{(3)}\cong\mathcal{B}_{<}^{(4)}$ to explain the bare essentials of the proof. We give a mathematically rigorous and complete proof of Eq. (8) in Appx. B for the interested readers.
@@ III.3 Isomorphisms between <$\mathcal{B}^{(k)}_{<}$’s
For $\alpha$ defined above, the, the following lemma holds, which is essential for the proof of Eq. (8).
@@
There is an isomorphism
$\mathcal{B}_{<}^{(4)}\ni\hat{Q}_{[4]}^{(4)}+\hat{Q}_{[4]}^{(3)}\mapsto\hat{Q}_{[3]}^{(3)}+\hat{Q}_{[3]}^{(2)}\in\mathcal{B}_{<}^{(3)},$ (32)
where $\hat{Q}_{[4]}^{(4)}$ and $\hat{Q}_{[3]}^{(3)}$ satisfy Eq. (24) for $k=4$.
@@
First we construct a map from $\mathcal{B}_{<}^{(4)}$ to $\mathcal{B}_{<}^{(3)}$. For $\hat{Q}_{[4]}^{(4)}\in\mathcal{B}_{\leq}^{(4)}$ and $\hat{Q}_{[4]}^{(3)}\in\mathcal{B}^{(3)}$, the condition $\hat{Q}_{[4]}^{(4)}+\hat{Q}_{[4]}^{(3)}\in\mathcal{B}_{<}^{(4)}$ is equivalent to the following equalities:
for all $i\in\Lambda_{N}$, where $\hat{Q}_{[2]}^{(2)}$ is defined by Eq. (28). Here, $[\ ,\hat{X}]:\mathcal{B}_{0}\ni\hat{Y}\mapsto[\hat{Y},\hat{X}]\in\mathcal{B}_{0}$ is the commutator with $\hat{X}\in\mathcal{B}_{0}$. Let us focus on the terms in the first bracket on the right-hand side (RHS) of Eq. (33). By applying Lem. 5 to the middle term, we can decompose these terms into the following form:
@@
Similarly to Prop. 4, there is an operator $\hat{Q^{\prime}}^{(2)}_{i+1}\in\mathcal{B}_{0}^{\otimes 2}$ satisfying
@@
for all $i\in\Lambda_{N}$, which is equivalent to $\hat{Q}_{[3]}^{(3)}+\hat{Q}_{[3]}^{(2)}\in\mathcal{B}_{<}^{(3)}$, where we define $\hat{Q}_{[3]}^{(2)}$ as $\frac{2}{3}\sum_{i=1}^{N}\hat{Q^{\prime}}^{(2)}_{i}$. Therefore, we have constructed a linear transformation $i_{34}:\mathcal{B}_{<}^{(4)}\to\mathcal{B}_{<}^{(3)}$. The injectivity of this map is confirmed from the injectivity (23). Conversely, by using the above equalities (e.g., Eq. (38)), we can construct a mapping $i_{43}:\mathcal{B}_{<}^{(3)}\to\mathcal{B}_{<}^{(4)}$ satisfying $i_{34}\circ i_{43}=\mathrm{id}_{\mathcal{B}_{<}^{(3)}}$. This implies the surjectivity of $i_{34}$, which completes the proof. ∎
III.4 Absence of spectrum generating algebras
As discussed in Ref. [25], the proof of the absence of local conserved quantities is applicable with a slight modification to the analysis of a related algebraic structure: spectrum generating algebras. A system with a Hamiltonian $\hat{H}$ is said to admit a spectrum generating algebra (SGA) for an operator $\hat{Q}$ if
@@ III.4 Absence of spectrum generating algebras
holds for some $\mathcal{E}\in\mathbb{R}\setminus\{0\}$. This algebraic structure is a variant of the condition $[\hat{Q},\hat{H}]=0$ for the existense of a conserved quantity, and $\eta$-pairing [59, 60] in the Hubbard model is a well-known example.
@@
Note that $\mathcal{B}_{<,\mathcal{E}}^{(k)}\subset\mathcal{B}_{\leq}^{(k)}\oplus\mathcal{B}^{(k-1)}$ holds. To prove Thm. 7 for the case of $k=4$, we first construct a linear transformation ${i^{\prime}}_{34}:\mathcal{B}_{<,\mathcal{E}}^{(4)}\to\mathcal{B}_{<,\frac{2}{3}\mathcal{E}}^{(3)}$.
In fact, this map can be defined in a manner similar to ${i}_{34}:\mathcal{B}_{<}^{(4)}\to\mathcal{B}_{<}^{(3)}$. For $\hat{Q}_{[4]}^{(4)}\in\mathcal{B}_{\leq}^{(4)}$ and $\hat{Q}_{[4]}^{(3)}\in\mathcal{B}^{(3)}$, the condition $\hat{Q}_{[4]}^{(4)}+\hat{Q}_{[4]}^{(3)}\in\mathcal{B}_{<,\mathcal{E}}^{(4)}$ is equivalent to the following equalities:
@@
$\hat{Q^{\prime}}^{(3)}_{L,i}\coloneqq\hat{Q}^{(3)}_{[4],i}+\hat{A}^{(3)}_{L,i}+\hat{B}^{(3)}_{L,i}-\mathcal{E}\hat{Q}^{(3)}_{[3],i}.$ (47)
@@
for all $i\in\Lambda_{N}$, which is equivalent to $\hat{Q}_{[3]}^{(3)}+\hat{Q}_{[3]}^{(2)}\in\mathcal{B}_{<,\frac{2}{3}\mathcal{E}}^{(3)}$.
On the other hand, for any $\mathcal{E}\neq 0$, $\hat{Q}_{[3]}^{(3)}+\hat{Q}_{[3]}^{(2)}\in\mathcal{B}_{<,\frac{2}{3}\mathcal{E}}^{(3)}$ implies $\hat{Q}_{[3]}^{(3)}=0$ by the assumption. Therefore, we have $\hat{Q}_{[4]}^{(4)}=0$, or equivalently, $\mathcal{B}_{<,\mathcal{E}}^{(4)}\subset\mathcal{B}^{(3)}$ for every $\mathcal{E}\in\mathbb{R}\setminus\{0\}$, which indicates that there is no $4$-local quantity satisfying Eq. (44). With a slight modification of the proof of Thm. 2 for the general $k$ case (Appx. B), we can show the general $k$ case as well. ∎
We note that SGAs are of interest in the study of thermalization. If an operator $\hat{Q}$ satisfies Eq. (44) on some subspaces, the system is said to admit a restricted spectrum generating algebra [61, 62]. This structure provides a unified origin of the tower of quantum many-body scar states, which prevents thermalization.
IV Algorithm of integrability test
By using Thm. 1 and Thm. 2, we can construct an algorithm of test for integrability and nonintegrability. In this section, we only consider translationally invariant Hamiltonians, for simplicity. In this case, we can decompose the set of $k$-local operators into momentum sectors by the Fourier transformation as
$\mathcal{B}^{(k)}=\bigoplus_{p\in\bigl{\{}e^{\frac{2\pi\mathrm{i}}{N}m}\bigm{|}m=0,\dots,N-1\bigr{\}}}\mathcal{B}^{(k,p)},$ (49)
where $p$ represents the momentum, which is the eigenvalue of the translation operation $\tau:\mathcal{B}^{\otimes A}\to\mathcal{B}^{\otimes(A-1)}$. Since this decomposition is consistent with the subspaces $\mathcal{B}^{(k)}_{\leq}$ and <$\mathcal{B}^{(k)}_{<}$, the integrability test introduced in Sec. II.2 is reduced to the following algorithm. First, we determine the space $\mathcal{B}^{(2,p)}_{\leq}\coloneqq\mathcal{B}^{(2)}_{\leq}\cap\mathcal{B}^{(2,p)}$. This step is equivalent to the following generalized eigenvalue problem:
@@ IV Algorithm of integrability test
for $\hat{Q}_{[2],1}^{(2)}\in\mathcal{B}_{0}^{\otimes 2}$, where we use the translational invariance $\tau(H^{(2)}_{i+1})=H^{(2)}_{i}$ and the fact that $\hat{Q}_{[2]}^{(2)}\in\mathcal{B}^{(2,p)}$ satisfies $\tau(\hat{Q}^{(2)}_{[2],i+1})=p\hat{Q}^{(2)}_{[2],i}$. We note that $\mathcal{B}^{(2,p)}_{\leq}=\{0\}$ except for at most $(d^{2}-1)^{2}$ numbers of ($N$-independent) momentum $p$ by the injectivity (23).

### v3

@@ IV.1 String Diagram
Before proving the theorems, we introduce a “basis-independent” approach for a proof of nonintegrability. We adopt a string diagram as a useful graphical language, which is widely used in various areas of physics and mathematics [59, 60, 61]. The basic idea is to represent a linear map using boxes and wires. For instance, $\hat{X}\in\mathcal{B}\otimes\mathcal{B}_{0}$, $f:\mathcal{B}_{0}^{\otimes 3}\to\mathcal{B}_{0}^{\otimes 2}$, and $(\mathrm{id}_{\mathcal{B}}\otimes f)(\hat{X}\otimes\bullet):\mathcal{B}_{0}^{\otimes 2}\to\mathcal{B}\otimes\mathcal{B}_{0}^{\otimes 2}$ are depicted as
@@
respectively, where $\mathrm{id}_{\mathcal{B}}$ is the identity map on $\mathcal{B}$. Here, each box represents a linear map: the top line corresponds to the input, and the bottom line to the output. In other words, time flows from top to bottom. Note that an element of a linear space is identified with a linear map from $\mathbb{C}$, which is represented by the “empty” line. The horizontal direction corresponds to the spatial direction, and placing the boxes horizontally corresponds to their tensor product. In the following, solid and dashed lines represent $\mathcal{B}_{0}$ and $\mathcal{B}$, respectively. Using this diagrammatic notation, the injectivity of $[\hat{H}^{(2)}_{i},\bullet\otimes I]$ and $[\hat{H}^{(2)}_{i},I\otimes\bullet]$ (see Eq. (4)) are equivalent to the existence of linear maps $\mu_{i},\lambda_{i}:\mathcal{B}_{0}^{\otimes 2}\to\mathcal{B}_{0}$ satisfying
@@
IV.2 Isomorphisms between $\mathcal{B}^{(k)}_{\leq}$’s
Now we move on to the proof of Thm. 2. It is sufficient to show the following proposition.
@@ IV.2 Isomorphisms between $\mathcal{B}^{(k)}_{\leq}$’s
for all $i\in\Lambda_{N}$, which is equivalent to the condition $\hat{Q}_{[k-1]}^{(k-1)}\in\mathcal{B}_{\leq}^{(k-1)}$. The uniqueness of $\hat{Q}_{[k-1]}^{(k-1)}$ follows from the injectivity (50). ∎
If we define $\hat{Q}_{[k]}^{(k)}\in\mathcal{B}^{(k)}$ as in Eq. (51) for some $\hat{Q}_{[k-1]}^{(k-1)}\in\mathcal{B}_{\leq}^{(k-1)}$, then it is easy to verify $\hat{Q}_{[k]}^{(k)}\in\mathcal{B}_{\leq}^{(k)}$. Therefore, we have established $\mathcal{B}_{\leq}^{(k)}\cong\mathcal{B}_{\leq}^{(k-1)}$ and Eq. (9). By applying Eq. (51) recursively, $\hat{Q}^{(k)}_{[k],i}$ is represented as
@@
where $\hat{Q}_{[2]}^{(2)}=\sum_{i=1}^{N}\hat{Q}^{(2)}_{[2],i}$ is an element of $\mathcal{B}^{(2)}_{\leq}$.
IV.3 Isomorphisms between <,$\mathcal{B}^{(k)}_{<,b}$’s
As pointed out at the end of Subsec. II.2, Thm. 1 follows from Thm. 3. Therefore, we conclude this section by proving Thm. 3. Here, we shall prove only the case of $\mathcal{B}_{<,b}^{(3)}\cong\mathcal{B}_{<,b}^{(4)}$ to explain the bare essentials of the proof. A rigorous and complete proof of Eq. (13) is provided in Appx. B for interested readers.
@@ IV.3 Isomorphisms between <,$\mathcal{B}^{(k)}_{<,b}$’s
For $\alpha$ defined above, the following lemma holds, which is essential for the proof of Eq. (13).
@@
There is a linear isomorphism
$\mathcal{B}_{<,b}^{(4)}\ni q\hat{Q}_{b}^{(4)}+\hat{Q}_{[4]}^{(3)}\mapsto q\hat{Q}_{b}^{(3)}+\hat{Q}_{[3]}^{(2)}\in\mathcal{B}_{<,b}^{(3)},$ (59)
where $q\in\mathbb{C}$ and $\hat{Q}_{b}^{(k)}$ is defined by Eq. (11).
@@
First we construct a map from $\mathcal{B}_{<,b}^{(4)}$ to $\mathcal{B}_{<,b}^{(3)}$. For $\hat{Q}_{[4]}^{(3)}\in\mathcal{B}^{(3)}$, the condition $q\hat{Q}_{b}^{(4)}+\hat{Q}_{[4]}^{(3)}\in\mathcal{B}_{<,b}^{(4)}$ is equivalent to the following equalities:
for all $i\in\Lambda_{N}$. Here, $[\ ,\hat{X}]:\mathcal{B}_{0}\ni\hat{Y}\mapsto[\hat{Y},\hat{X}]\in\mathcal{B}_{0}$ is the commutator with $\hat{X}\in\mathcal{B}_{0}$. Let us focus on the terms in the first bracket on the right-hand side (RHS) of Eq. (60). By applying Lem. 7 to the middle term, we can decompose these terms into the following form:
@@
Similarly to Prop. 6, there is an operator $\hat{Q^{\prime}}^{(2)}_{i+1}\in\mathcal{B}_{0}^{\otimes 2}$ satisfying
@@
for all $i\in\Lambda_{N}$, which is equivalent to $q\hat{Q}_{b}^{(3)}+\hat{Q}_{[3]}^{(2)}\in\mathcal{B}_{<,b}^{(3)}$, where we define $\hat{Q}_{[3]}^{(2)}$ as $\frac{2}{3}\sum_{i=1}^{N}\hat{Q^{\prime}}^{(2)}_{i}$. Therefore, we have constructed a linear transformation $i_{34}:\mathcal{B}_{<,b}^{(4)}\to\mathcal{B}_{<,b}^{(3)}$. The injectivity of this map is confirmed from the injectivity (50). Conversely, by using the above equalities (e.g., Eq. (65)), we can construct a mapping $i_{43}:\mathcal{B}_{<}^{(3)}\to\mathcal{B}_{<}^{(4)}$ satisfying $i_{34}\circ i_{43}=\mathrm{id}_{\mathcal{B}_{<}^{(3)}}$. This implies the surjectivity of $i_{34}$. Furthermore, the action of $i_{43}$ on $\mathcal{B}_{\leq}^{(2)}$ is the same as that of $\frac{2}{3}\iota_{2}$, which completes the proof. ∎
IV.4 Absence of spectrum generating algebras
As discussed in Ref. [25], the proof for the absence of local conserved quantities can be slightly modified to apply to the analysis of a related algebraic structure: spectrum generating algebras. A system with a Hamiltonian $\hat{H}$ is said to admit a spectrum generating algebra (SGA) for an operator $\hat{Q}$ if
@@ IV.4 Absence of spectrum generating algebras
holds for some $\mathcal{E}\in\mathbb{R}\setminus\{0\}$. This algebraic structure is a variant of the condition $[\hat{Q},\hat{H}]=0$ for the existence of a conserved quantity, and $\eta$-pairing [63, 64] in the Hubbard model is a well-known example.
@@
Note that $\mathcal{B}_{<,\mathcal{E}}^{(k)}\subset\mathcal{B}_{\leq}^{(k)}\oplus\mathcal{B}^{(k-1)}=\mathbb{C}\hat{Q}_{b}^{(k)}\oplus\mathcal{B}^{(k-1)}$ holds. To prove Thm. 9 for the case of $k=4$, we first construct a linear transformation ${i^{\prime}}_{34}:\mathcal{B}_{<,\mathcal{E}}^{(4)}\to\mathcal{B}_{<,\frac{2}{3}\mathcal{E}}^{(3)}$.
In fact, this map can be defined in a manner similar to ${i}_{34}:\mathcal{B}_{<}^{(4)}\to\mathcal{B}_{<}^{(3)}$. For $q\in\mathbb{C}$ and $\hat{Q}_{[4]}^{(3)}\in\mathcal{B}^{(3)}$, the condition $q\hat{Q}_{b}^{(4)}+\hat{Q}_{[4]}^{(3)}\in\mathcal{B}_{<,\mathcal{E}}^{(4)}$ is equivalent to the following equalities:
@@
$\hat{Q^{\prime}}^{(3)}_{L,i}\coloneqq\hat{Q}^{(3)}_{[4],i}+\hat{A}^{(3)}_{L,i}+\hat{B}^{(3)}_{L,i}-\mathcal{E}q\hat{Q}_{b,i}^{(3)}.$ (74)
@@
for all $i\in\Lambda_{N}$, which is equivalent to $q\hat{Q}_{b}^{(3)}+\hat{Q}_{[3]}^{(2)}\in\mathcal{B}_{<,\frac{2}{3}\mathcal{E}}^{(3)}$.
On the other hand, for any $\mathcal{E}\neq 0$, $\hat{Q}_{[3]}^{(3)}+\hat{Q}_{[3]}^{(2)}\in\mathcal{B}_{<,\frac{2}{3}\mathcal{E}}^{(3)}$ implies $q=0$ by the assumption. Therefore, we have $\mathcal{B}_{<,\mathcal{E}}^{(4)}\subset\mathcal{B}^{(3)}$ for every $\mathcal{E}\in\mathbb{R}\setminus\{0\}$, which indicates that there is no $4$-local quantity satisfying Eq. (71). With a slight modification of the proof of Thm. 3 for the general $k$ case (Appx. B), we can show the general $k$ case as well. ∎
We note that SGAs are of interest in the study of thermalization. If an operator $\hat{Q}$ satisfies Eq. (71) on some subspaces, the system is said to admit a restricted spectrum generating algebra [65, 66]. This structure offers a unified origin of the tower of quantum many-body scar states, which prevents thermalization.
V Algorithm of integrability test
Using Thm. 1 and Thm. 2, we can construct an algorithm to test the integrability and nonintegrability of one-dimensional quantum spin systems. In this section, we only consider translationally invariant systems. In this case, the set of $k$-local operators can be decomposed into momentum sectors via the Fourier transformation as
$\mathcal{B}^{(k)}=\bigoplus_{p\in\bigl{\{}\frac{2\pi}{N}m\bigm{|}m=0,\dots,N-1\bigr{\}}}\mathcal{B}^{(k,p)},$ (76)
where $p$ denotes the momentum, which corresponds to the eigenvalue of the translation operation $\tau:\mathcal{B}^{\otimes A}\to\mathcal{B}^{\otimes(A-1)}$. Since this decomposition is compatible with the subspaces $\mathcal{B}^{(k)}_{\leq}$ and <$\mathcal{B}^{(k)}_{<}$, the integrability test introduced in Sec. II.2 can be reduced to the following algorithm. First, we determine the subspace $\mathcal{B}^{(2,p)}_{\leq}\coloneqq\mathcal{B}^{(2)}_{\leq}\cap\mathcal{B}^{(2,p)}$, which is equivalent to solving the following generalized eigenvalue problem:
@@ V Algorithm of integrability test
for $\hat{Q}_{[2],1}^{(2)}\in\mathcal{B}_{0}^{\otimes 2}$, where we use the translational invariance $\tau(H^{(2)}_{i+1})=H^{(2)}_{i}$ and the fact that $\hat{Q}_{[2]}^{(2)}\in\mathcal{B}^{(2,p)}$ satisfies $\tau(\hat{Q}^{(2)}_{[2],i+1})=e^{\mathrm{i}p}\hat{Q}^{(2)}_{[2],i}$. We note that $\mathcal{B}^{(2,p)}_{\leq}=\{0\}$ except for at most $(d^{2}-1)^{2}$ numbers of ($N$-independent) momentum $p$ due to the injectivity (50). If $\mathcal{B}_{\leq}^{(2,p=0)}=\mathbb{C}\hat{H}^{(2)}$, we can move on to the next step.